OpenStudy (anonymous):

Simplify sin(45-x)cos(45-x)

hero (hero):

Used Sum and difference Angle formulas for sine and cosine to simplify

OpenStudy (anonymous):

So 2sinAcosB=sin(A+B)+sin(A-B)

OpenStudy (anonymous):

But how the question sin and cos become sin and sin in the forumla?

hero (hero):

where'd you get that formula.

OpenStudy (anonymous):

My tudor

OpenStudy (anonymous):

Is it incorrect?

hero (hero):

It doesn't look correct as written

OpenStudy (anonymous):

What's the correct formula?

OpenStudy (anonymous):

Ohh, right

OpenStudy (anonymous):

So what formula do I use?

OpenStudy (anonymous):

Do I work with sin (45-x) separately and then cos (45-x) separately?

OpenStudy (anonymous):

Yep, I understood that part

OpenStudy (anonymous):

I subbed in the values and got this:

OpenStudy (anonymous):

(sin45cosx-cos45sinx) (cos45cosx+sin45sinx)

hero (hero):

You should know what Sin 45 and cos ( 45)

OpenStudy (anonymous):

Sin 45 is rt2/2

OpenStudy (anonymous):

And cos 45 is rt2/3

hero (hero):

They're both sqrt(2)/2

OpenStudy (anonymous):

What about the x ?

hero (hero):

You don't need to Know What x ls. You're simplifying an expression not solving an equation.

OpenStudy (anonymous):

Am I doing this wrong?

hero (hero):

Yes

hero (hero):

Okay

OpenStudy (anonymous):

How do I further simplify this?

hero (hero):

Algebraically

OpenStudy (anonymous):

So that becomes

OpenStudy (anonymous):

(rt2/2)^2 x (cosx-sinx)(cosx+sinx)

hero (hero):

Just try to work on it and post your simplification

OpenStudy (anonymous):

= 1/2 x (cosx-sinx)(cosx+sinx)

OpenStudy (anonymous):

I don't know how to further simplify

hero (hero):

There's no x after 1/2

OpenStudy (anonymous):

Oh that's a multiplication sign

OpenStudy (anonymous):

= 1/2(cosx-sinx)(cosx+sinx)

OpenStudy (anonymous):

That's a difference of two squares

hero (hero):

correct

OpenStudy (anonymous):

=1/2(cosx-sinx)^2

hero (hero):

Not written correctly

hero (hero):

Do you want to try to correct your own Mistake?

OpenStudy (anonymous):

Is it 1/2 cosx^2-sin^2 ?

hero (hero):

That's better

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Ok, how do I further simplify?

hero (hero):

sin(×) is the primitive trig expression So express the simplification in terms Of sine.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What do you mean?

hero (hero):

Do you know the Pythagorean identity?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sin^2x+cos^2x=1

OpenStudy (anonymous):

Could you please give me the completed working out, I'm a bit short on time

hero (hero):

Isolate cos^2 x

hero (hero):

We're almost done.

OpenStudy (anonymous):

And then..?

hero (hero):

substitute cos ^2x with its equivalent expression then finish simplifying

OpenStudy (anonymous):

Ok thank you so much

hero (hero):

By the way,

OpenStudy (anonymous):

Yes?

hero (hero):